Question 8 3.35/5 Kansas Company uses a standard cost accounting system. In 2020, the company produced 28,500 units. Each unit took several pounds of direct materials and 1.6 standard hours of direct labor at a standard hourly rate of $11.00. Normal capacity was 50,000 direct labor hours. During the year, 131,700 pounds of raw materials were purchased at $0.95 per pound. All materials purchased were used during the year. (a) Correct answer iconYour answer is correct. If the materials price variance was $5,268 favorable, what was the standard materials price per pound? (Round answer to 2 decimal places, e.g. 2.75.) Standard materials price per pound $ Attempts: 1 of 4 used (b) Correct answer iconYour answer is correct. If the materials quantity variance was $17,523 unfavorable, what was the standard materials quantity per unit? Standard materials quantity per unit pounds Attempts: 1 of 4 used (c) Correct answer iconYour answer is correct. What were the standard hours allowed for the units produced? Standard hours allowed hours Attempts: 1 of 4 used (d) Correct answer iconYour answer is correct. If the labor quantity variance was $6,600 unfavorable, what were the actual direct labor hours worked? Actual hours worked hours Attempts: 1 of 4 used (e) Correct answer iconYour answer is correct. If the labor price variance was $4,620 favorable, what was the actual rate per hour? (Round answer to 2 decimal places, e.g. 2.75.) Actual rate per hour $ Attempts: 1 of 4 used (f) Correct answer iconYour answer is correct. If total budgeted manufacturing overhead was $325,000 at normal capacity, what was the predetermined overhead rate based on direct labor hours? (Round answer to 2 decimal places, e.g. 2.75.) Predetermined overhead rate $ Attempts: 1 of 4 used (g) Incorrect answer iconYour answer is incorrect. What was the standard cost per unit of product? (Round answer to 2 decimal places, e.g. 2.75.) Standard cost per unit $ eTextbook and Media Attempts: 2 of 4 used Save for LaterCheck Answer
Answer #1
a) Material price variance=Actual material used*(Standard rate for material-Actual rate for material)
Material price variance=$ 5268
Actual material used=131700 pounds
Standard rate for material=SR
Actual rate=45600/12000=$0.95 per pound
5268=131700*(SR-0.95)
SR-0.95=5268/131700
SR-0.95=0.04
SR=0.04+095=$0.99 per pound
b) Material quantity variance=Standard rate for materials*(Standard materials required-Actual material used)
Material quantity variance=-17523 (Unfavorable)
Standard rate=$ 0.99 per pound
Standard materials required=SQ
Actual material used=131700 pounds
(17523)=0.99*(SQ-131700)
SQ-131700=-17523/0.99
SQ-131700=-17700
SQ=-17700+131700=$114000
Standard materials per unit=SQ/Units produced=114000/28500=4 pounds per unit
c) Standard hours allowed for the units produced=Units produced*Standard hours required per unit=28500*1.6=45600 hours
d) Labor quantity variance=Standard labor rate*(Standard labor hours required-Actual labor hours worked)
Labor quantity variance=-6600 (Unfavorable)
Standard labor rate=$11 per hour
Standard labor hours required=45600 hours
Actual hours worked=AH
(6600)=11*(45600-AH)
45600-AH=-6600/11
45600-AH=-600
AH=45600+600=46200 hours
e) Labor price variance=Actual hours worked*(Standard labor rate-Actual labor rate)
Labor price variance=4620
Actual hours worked=46200 hours
Standard labor rate=$11 per hour
Actual labor rate=AR
4620=46200*(11-AR)
11-AR=4620/46200
11-AR=0.10
AR=11-0.10=$10.9 per hour
f) Predetermined overhead rate=Budgeted manufacturing overhead at normal capacity/direct labor hours at normal capacity=325000/50000=$6.5 per labor hour
g) Standard cost per unit:
$
Direct material
(Std qty*Std rate) (4*0.99) 3.96
Direct labor
(Std hrs*std rate) (1.6*11) 17.6
Manufacturing overhead
(1.6*6.5) 10.4
Total 31.96